# StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions

### v2

@@ Abstract
Visual Analytics for Education, Student-ChatGPT Interaction
@@
System interface of StuGPTViz. The Filter View (A) offers an overview and enables filtering of tasks and students through the Task Overview (a1) and Student Overview (a2). The Pattern View (B) displays the macro-level summary of conversation characteristics in the Pattern Summary (b1), and micro-level interaction patterns along with their evolution in the Pattern Nuance, represented by the Pattern Mining Tabl (b4) and the Interaction Tree (b5). The Detail View (C) presents task descriptions (c1), and the raw conversation data between students and ChatGPT (c2).
@@ 1.3 Visual Analytics for Conversational Data
The exploration of visual analytics for conversational text data within the visualization community has encompassed a wide range of applications, from sentiment analysis and topic modeling to mapping conversation flows and interactions within user groups [36, 46, 30, 32]. For instance, T-Cal and IneqDetect [31, 54] focus on analyzing group conversations and estimating members’ sentiments to assess collaboration effectiveness. Meanwhile, efforts such as ThreadReconstructor, MultiConVis, and VisOHC [24, 37, 6] probe into the structure and core topics of online forum discussions. However, these initiatives mainly focus on summarizing the dynamics of multi-party conversations without delving into the intricacies of one-on-one dialogues.
@@ 3 Design Requirement and Data Processing
This section presents the design requirements for the visual analytics system identified through discussions with our experts (E1 and E2), the procedure for coding student prompts, and the methodology for processing ChatGPT responses.
@@ 4.3.1 Pattern Summary
The Pattern Summary is introduced with a control button at the top-left corner (Fig. 3) for selecting the grouping mode. A comprehensive macro-level pattern summary (R2) under the chosen mode is displayed below. The grey card component presents a between-group level summary (Fig. 3-A), adjacent to which are the within-group level summary cards (Fig. 3-B). To begin, instructors can view the summary donuts chart (Fig. 3-a1), which shows the distribution of all students’ prompt categories as defined in Section 3.2. For example, here, a predominantly light yellow section suggests that a majority of student prompts are at the “Remember” cognitive stage, while minor blue and major green segments indicate a limited use of literature-supported effective prompt strategies. This donuts chart design aims to clearly display the percentage of each cognitive level in students’ prompts while optimizing space usage. Additionally, to effectively display and compare the overall quality of ChatGPT responses and students’ learning outcomes, we selected two light grey bars and a dark grey bar (Fig. 3-a1) to represent the three metrics due to their simplicity and effectiveness [56]. The details of metrics are introduced in Section 3.3.
Instructors can easily check each group’s background information by hovering over the summary donut chart (Fig. 4-A) and perform between-group comparisons of cognitive stage distribution and ChatGPT response quality using stacked bar charts and accompanying grey bar charts (Fig. 3-a2). Moreover, by switching to “TaskG” mode (Fig. 4-B), instructors can change the grouping from student background to task types, shifting the analysis towards task-specific student performance and interaction summaries. This flexible approach enables an efficient multi-viewpoint summary of macro-level students-ChatGPT conversations (R2). After identifying a specific group of interest, instructors can click on a stacked bar in the summary card (Fig. 3-a2), which highlights the selected group card (Fig. 3-b1) for deeper, within-group analysis. Instructors can sort students by various metrics by clicking on the metric summary bars in the first row, as shown in the example where students are sorted by their scores in ascending order (Fig. 3-b1). To move on, instructors can click on any stacked bar or the donuts chart (Fig. 3-b1) to investigate micro-level details in Pattern Mining Table (Fig. 5-A) and the Interaction Tree (Fig. 6-A).
@@ 4.3.3 Interaction Tree
The Interaction Tree employs a decision-tree format design to trace the evolution of interaction patterns (R4) for individual students under each task (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions, Fig. 6-A). Each path within the Interaction Tree represents a student’s detailed interaction process with ChatGPT under a specific task, beginning from a common root where nodes symbolize students’ prompts, and adjacent solid links depict ChatGPT’s responses. This format enables the aggregation of similar interaction paths, highlighting variations and commonalities in student interaction patterns. Consistent with the above settings, colors in the node coding for the category of the prompt and abbreviations inside or beside the node detailing the code content. When prompts encompass multiple codes, nodes incorporate pie-chart coloring to represent each code visually (Fig. 6-a1). To aggregate prompts with the same code contents, we use node size to indicate the number of students at the same round in their conversation. On the other hand, the solid links between nodes convey ChatGPT’s response characteristics, including “Response Token Length” (RL) and “Information Gain” (IG). Each ChatGPT response’s “Information Gain” is represented through variations in the horizontal length of the link, as it is the major information instructors (E1 & E2) want to notice, while link width and opacity double encode “Response Token Length”. This visual encoding offers insights into the value and quality of ChatGPT’s replies. Furthermore, the end of each path features one grey tag representing the student’s alias and task’s ID (Fig. 6-a2), together with another tag below encoding the student’s performance on this task. Here, we utilized numerical values and color intensity to denote scores, thereby linking interaction patterns directly to learning outcomes. For instance, the highlighted student gained the full mark (Fig. 6-a2) after various interactions with ChatGPT. Through these meticulously designed components, instructors are not only equipped to perform a granular analysis of student interactions but also able to correlate them with educational outcomes, facilitating a comprehensive understanding of students’ learning behaviors and the effectiveness of their interactions with ChatGPT (R5).
@@ 5.3.1 Case 1: Enhancing Students’ ChatGPT Utilization
Figure 7: (A) The interaction pattern summary within Group 1. (B) The pattern table, sorting all interaction patterns mined from students with alias “cs”. (C) The descriptions for T14 & T6 and the student’s key prompts. (D) The pattern paths of student “cs” for T14 and T6.
@@
Analysis of Interaction Patterns: The examination of “Task 3” through the pattern mining table and Interaction Tree (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b4, b5) revealed nuanced dynamics of student-ChatGPT interactions. Initially, sorting by average score revealed infrequent and overly specific patterns. To uncover broadly applicable patterns, experts shifted to sorting by frequency (“Count”), identifying a recurring and effective pattern: a combination of “Definition Inquiry” and “Follow up Questions” (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b6), notable for its prevalence and high average score exceeding 0.8. Experts agreed this simple yet effective combination was crucial for foundational understanding and deeper exploration of ChatGPT’s responses. To observe students using this pattern, experts selected it in the table, highlighting corresponding students in the Interaction Tree (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b8, Fig. 6-B). A standout instance involved a student, “CJ”, who used this pattern to achieve a full score with minimal conversation (Fig. 6-b2). By clicking the student-task ID tag (Fig. 6-b2) and reviewing the raw conversations (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-c3) along the tree branch, experts saw “CJ” begin with a “Definition Inquiry”, asking if ChatGPT understood the pipeline concept [61]. After confirmation, “CJ” asked a “Follow up Question” for further exploration. Noticing that ChatGPT did not provide answers from the multiple choices listed in the task description, “Output Restrictions” were strategically implemented. Despite these restrictions, ChatGPT’s responses remained flawed, leading “CJ” to revert to “Definition Inquiry” for double verification before progressing. Once a correct response was secured from ChatGPT, “CJ” advanced through the remaining sections, receiving partially correct responses from ChatGPT but ultimately providing completely correct answers independently. Experts concluded that “Output Restrictions” in prompts maximize ChatGPT’s utility for precise information retrieval. Revisiting the pattern mining table (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b4), they noted “Output Restrictions” as a significant pattern (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b7) due to its high frequency and average score. Thus, the experts finally identified the combination of “Definition Inquiry” and “Follow up Questions” with “Output Restrictions” as a potent strategy for engaging with ChatGPT effectively. The raw conversation data of “CJ” is provided in the supplementary (C). This evaluation also highlighted ChatGPT’s limitations in addressing abstract questions, advising against outright reliance on its answers but encouraging a focus on its reasoning and factual accuracy. Ultimately, the experts planned to incorporate these insights into feedback for their students.
@@ 6 Discussion
ChatGPT for Data Visualization Education The introduction of ChatGPT into educational ecosystems marks a pivotal moment and necessitates a nuanced understanding of technological integration in learning. Our investigation into the patterns and strategies of student engagement with ChatGPT is critical, providing insights that guide instructors in facilitating the effective use of AI. Our study revealed several key findings from student questionnaires, student-ChatGPT conversation data, and discussions with course instructors. First, students reported that ChatGPT excels in summarizing key concepts, providing quick access to vast information, and offering tailored Q&A sessions. Instructors agreed that these functionalities are particularly beneficial for students with limited backgrounds, such as those who changed their major in graduate school, allowing them to keep pace with coursework without hindering class progress. Additionally, over $90\%$ of students expressed satisfaction with ChatGPT’s ability to handle data visualization queries, indicating a strong positive perception of its utility in data visualization education. However, instructors pointed out that recognizing ChatGPT’s limitations in interpreting and processing visual data compared to textual information is crucial. For instance, through the collected student-ChatGPT conversations, instructors identified that some students resorted to asking ChatGPT for URLs of existing sample visualizations (e.g., demos of parallel coordinates) to obtain better figure quality after receiving a low-quality response. These findings inspire instructors to develop innovative approaches, such as guiding ChatGPT to provide descriptions or links to reference visualizations, for better educational content. Instructors also emphasized that the demand-driven nature of data visualization requires students to employ precise and strategic questioning techniques. This underscores the importance of teaching students how to prompt effectively to maximize ChatGPT’s capabilities. Furthermore, as we navigate the ChatGPT-enhanced educational paradigm, our work highlighted the need to focus on cultivating students’ higher-order cognitive skills, such as critical thinking and evaluative judgment. For instance, instructors suggested encouraging students to ask ChatGPT for alternative designs with underlying rationales, enabling them to critically assess options and fostering a collaborative learning dynamic. Approaches like this redefine the role of ChatGPT in education—from a universal solver to a pedagogical partner—and emphasize the importance of critical engagement and decision-making skills in the data visualization domain.
@@
Generalizability & Scalability The application of StuGPTViz, while rooted in the context of data visualization education, unveils broader implications for ChatGPT-assisted learning environments. It offers a robust framework for analyzing student interactions with ChatGPT across a variety of courses. StuGPTViz’s core workflow, which includes filtering tasks and students of interest, analyzing cognitive levels through prompts, assessing prompt engineering skills, evaluating ChatGPT’s response quality, and identifying interaction patterns, encapsulates the universal aspects of leveraging ChatGPT in education. This approach enriches our understanding of effective AI integration into pedagogy and opens avenues for examining ethical considerations regarding ChatGPT’s involvement in teaching practices. Regarding scalability, StuGPTViz demonstrates competence in managing classes of 48 students through the interaction tree visualization techniques equipped with simple pruning for clarity. This capability ensures the StuGPTViz’s efficacy in distilling actionable insights from complex datasets, catering to the needs of regular-sized classes. However, scalability challenges arise as class sizes expand beyond this scope. For larger cohorts (e.g., exceeding 100 students) the incorporation of edge bundling techniques emerges as a potential refinement to enhance pattern visualization and comparison. Additionally, expanding the pattern mining table to include interaction patterns from different tasks or courses would further enhance its scalability and provide additional benefits. This adaptation will form a component of our further efforts, aiming to ensure that StuGPTViz remains effective in in diverse educational settings, advancing the goal of inclusive AI-enhanced education.

### v3

@@ Abstract
Visual analytics for education, ChatGPT for education, student-ChatGPT interaction
@@
System interface of StuGPTViz. The Filter View (A) offers an overview and enables filtering of tasks and students through the Task Overview (a1) and Student Overview (a2). The Pattern View (B) displays the macro-level summary of conversation characteristics in the Pattern Summary (b1), and micro-level interaction patterns along with their evolution in the Pattern Nuance, represented by the Pattern Mining Table (b4) and the Interaction Tree (b5). The Detail View (C) presents task descriptions (c1), and the raw conversation data between students and ChatGPT (c2).
@@ 1.3 Visual Analytics for Conversational Data
The exploration of visual analytics for conversational text data within the visualization community has encompassed a wide range of applications, from sentiment analysis and topic modeling to mapping conversation flows and interactions within user groups [35, 45, 31]. For instance, T-Cal and IneqDetect [30, 53] focus on analyzing group conversations and estimating members’ sentiments to assess collaboration effectiveness. Meanwhile, efforts such as ThreadReconstructor, MultiConVis, and VisOHC [24, 36, 6] probe into the structure and core topics of online forum discussions. However, these initiatives mainly focus on summarizing the dynamics of multi-party conversations without delving into the intricacies of one-on-one dialogues.
@@ 3 Design Requirement and Data Processing
This section presents the design requirements for the visual analytics system identified through discussions with our experts (E1 & E2), the procedure for coding student prompts, and the methodology for processing ChatGPT responses.
@@ 4.3.1 Pattern Summary
The Pattern Summary is introduced with a control button at the top-left corner (Fig. 3) for selecting the grouping mode. A comprehensive macro-level pattern summary (R2) under the chosen mode is displayed below. The grey card component presents a between-group level summary (Fig. 3-A), adjacent to which are the within-group level summary cards (Fig. 3-B). To begin, instructors can view the summary donut chart (Fig. 3-a1), which shows the distribution of all students’ prompt categories as defined in Section 3.2. For example, here, a predominantly light yellow section suggests that a majority of student prompts are at the “Remember” cognitive stage, while minor blue and major green segments indicate a limited use of literature-supported effective prompt strategies. This donut chart design aims to clearly display the percentage of each cognitive level in students’ prompts while optimizing space usage. Additionally, to effectively display and compare the overall quality of ChatGPT responses and students’ learning outcomes, we selected two light grey bars and a dark grey bar (Fig. 3-a1) to represent the three metrics due to their simplicity and effectiveness [55]. The details of metrics are introduced in Section 3.3.
Instructors can easily check each group’s background information by hovering over the summary donut chart (Fig. 4-A) and perform between-group comparisons of cognitive stage distribution and ChatGPT response quality using stacked bar charts and accompanying grey bar charts (Fig. 3-a2). Moreover, by switching to “TaskG” mode (Fig. 4-B), instructors can change the grouping from student background to task types, shifting the analysis towards task-specific student performance and interaction summaries. This flexible approach enables an efficient multi-viewpoint summary of macro-level students-ChatGPT conversations (R2). After identifying a specific group of interest, instructors can click on a stacked bar in the summary card (Fig. 3-a2), which highlights the selected group card (Fig. 3-b1) for deeper, within-group analysis. Instructors can sort students by various metrics by clicking on the metric summary bars in the first row, as shown in the example where students are sorted by their scores in ascending order (Fig. 3-b1). To move on, instructors can click on any stacked bar or the donut chart (Fig. 3-b1) to investigate micro-level details in Pattern Mining Table (Fig. 5-A) and the Interaction Tree (Fig. 6-A).
@@ 4.3.3 Interaction Tree
The Interaction Tree employs a decision-tree format design to trace the evolution of interaction patterns (R4) for individual students under each task (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b5, Fig. 6-A). Each path within the Interaction Tree represents a student’s detailed interaction process with ChatGPT under a specific task, beginning from a common root where nodes symbolize students’ prompts, and adjacent solid links depict ChatGPT’s responses. This format enables the aggregation of similar interaction paths, highlighting variations and commonalities in student interaction patterns. Consistent with the above settings, colors in the node coding for the category of the prompt and abbreviations inside or beside the node detailing the code content. When prompts encompass multiple codes, nodes incorporate pie-chart coloring to represent each code visually (Fig. 6-a1). To aggregate prompts with the same code contents, we use node size to indicate the number of students at the same round in their conversation. On the other hand, the solid links between nodes convey ChatGPT’s response characteristics, including “Response Token Length” (RL) and “Information Gain” (IG). Each ChatGPT response’s “Information Gain” is represented through variations in the horizontal length of the link, as it is the major information instructors (E1 & E2) want to notice, while link width and opacity double encode “Response Token Length”. This visual encoding offers insights into the value and quality of ChatGPT’s replies. Furthermore, the end of each path features one grey tag representing the student’s alias and task’s ID (Fig. 6-a2), together with another tag below encoding the student’s performance on this task. Here, we utilized numerical values and color intensity to denote scores, thereby linking interaction patterns directly to learning outcomes. For instance, the highlighted student gained the full mark (Fig. 6-a2) after various interactions with ChatGPT. Through these meticulously designed components, instructors are not only equipped to perform a granular analysis of student interactions but also able to correlate them with educational outcomes, facilitating a comprehensive understanding of students’ learning behaviors and the effectiveness of their interactions with ChatGPT (R5).
@@ 5.3.1 Case 1: Enhancing Students’ ChatGPT Utilization
Figure 7: (A) The interaction pattern summary within Group 1. (B) The pattern table, sorting all interaction patterns mined from students with alias “cx”. (C) The descriptions for T14 & T6 and the student’s key prompts. (D) The pattern paths of student “cx” for T14 and T6.
@@
Analysis of Interaction Patterns: The examination of “Task 3” through the pattern mining table and Interaction Tree (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b4, b5) revealed nuanced dynamics of student-ChatGPT interactions. Initially, sorting by average score revealed infrequent and overly specific patterns. To uncover broadly applicable patterns, experts shifted to sorting by frequency (“Count”), identifying a recurring and effective pattern: a combination of “Definition Inquiry” and “Follow up Questions” (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b6), notable for its prevalence and high average score exceeding 0.8. Experts agreed this simple yet effective combination was crucial for foundational understanding and deeper exploration of ChatGPT’s responses. To observe students using this pattern, experts selected it in the table, highlighting corresponding students in the Interaction Tree (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b8, Fig. 6-A). A standout instance involved a student, “CJ”, who used this pattern to achieve a full score with minimal conversation (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b8). By clicking the student-task ID tag (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b8) and reviewing the raw conversations (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-c2) along the tree branch, experts saw “CJ” begin with a “Definition Inquiry”, asking if ChatGPT understood the pipeline concept [60]. After confirmation, “CJ” asked a “Follow up Question” for further exploration. Noticing that ChatGPT did not provide answers from the multiple choices listed in the task description, “Output Restrictions” were strategically implemented. Despite these restrictions, ChatGPT’s responses remained flawed, leading “CJ” to revert to “Definition Inquiry” for double verification before progressing. Once a correct response was secured from ChatGPT, “CJ” advanced through the remaining sections, receiving partially correct responses from ChatGPT but ultimately providing completely correct answers independently. Experts concluded that “Output Restrictions” in prompts maximize ChatGPT’s utility for precise information retrieval. Revisiting the pattern mining table (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b4), they noted “Output Restrictions” as a significant pattern (StuGPTViz: A Visual Analytics Approach to Understand Student-ChatGPT Interactions-b7) due to its high frequency and average score. Thus, the experts finally identified the combination of “Definition Inquiry” and “Follow up Questions” with “Output Restrictions” as a potent strategy for engaging with ChatGPT effectively. The raw conversation data of “CJ” is provided in the supplementary (C). This evaluation also highlighted ChatGPT’s limitations in addressing abstract questions, advising against outright reliance on its answers but encouraging a focus on its reasoning and factual accuracy. Ultimately, the experts planned to incorporate these insights into feedback for their students.
@@ 6 Discussion
ChatGPT for Data Visualization Education: The introduction of ChatGPT into educational ecosystems marks a pivotal moment and necessitates a nuanced understanding of technological integration in learning. Our investigation into the patterns and strategies of student engagement with ChatGPT is critical, providing insights that guide instructors in facilitating the effective use of AI. Our study revealed several key findings from student questionnaires, student-ChatGPT conversation data, and discussions with course instructors. First, students reported that ChatGPT excels in summarizing key concepts, providing quick access to vast information, and offering tailored Q&A sessions. Instructors agreed that these functionalities are particularly beneficial for students with limited backgrounds, such as those who changed their major in graduate school, allowing them to keep pace with coursework without hindering class progress. Additionally, over $90\%$ of students expressed satisfaction with ChatGPT’s ability to handle data visualization queries, indicating a strong positive perception of its utility in data visualization education. However, instructors pointed out that recognizing ChatGPT’s limitations in interpreting and processing visual data compared to textual information is crucial. For instance, through the collected student-ChatGPT conversations, instructors identified that some students resorted to asking ChatGPT for URLs of existing sample visualizations (e.g., demos of parallel coordinates) to obtain better figure quality after receiving a low-quality response. These findings inspire instructors to develop innovative approaches, such as guiding ChatGPT to provide descriptions or links to reference visualizations, for better educational content. Instructors also emphasized that the demand-driven nature of data visualization requires students to employ precise and strategic questioning techniques. This underscores the importance of teaching students how to prompt effectively to maximize ChatGPT’s capabilities. Furthermore, as we navigate the ChatGPT-enhanced educational paradigm, our work highlighted the need to focus on cultivating students’ higher-order cognitive skills, such as critical thinking and evaluative judgment. For instance, instructors suggested encouraging students to ask ChatGPT for alternative designs with underlying rationales, enabling them to critically assess options and fostering a collaborative learning dynamic. Approaches like this redefine the role of ChatGPT in education—from a universal solver to a pedagogical partner—and emphasize the importance of critical engagement and decision-making skills in the data visualization domain.
@@
Generalizability & Scalability: The application of StuGPTViz, while rooted in the context of data visualization education, unveils broader implications for ChatGPT-assisted learning environments. It offers a robust framework for analyzing student interactions with ChatGPT across a variety of courses. StuGPTViz’s core workflow, which includes filtering tasks and students of interest, analyzing cognitive levels through prompts, assessing prompt engineering skills, evaluating ChatGPT’s response quality, and identifying interaction patterns, encapsulates the universal aspects of leveraging ChatGPT in education. This approach enriches our understanding of effective AI integration into pedagogy and opens avenues for examining ethical considerations regarding ChatGPT’s involvement in teaching practices. Regarding scalability, StuGPTViz demonstrates competence in managing classes of 48 students through the interaction tree visualization techniques equipped with simple pruning for clarity. This capability ensures the StuGPTViz’s efficacy in distilling actionable insights from complex datasets, catering to the needs of regular-sized classes. However, scalability challenges arise as class sizes expand beyond this scope. For larger cohorts (e.g., exceeding 100 students) the incorporation of edge bundling techniques emerges as a potential refinement to enhance pattern visualization and comparison. Additionally, expanding the pattern mining table to include interaction patterns from different tasks or courses would further enhance its scalability and provide additional benefits. This adaptation will form a component of our further efforts, aiming to ensure that StuGPTViz remains effective in in diverse educational settings, advancing the goal of inclusive AI-enhanced education.
@@ 7 Conclusion
Acknowledgements.
This work has been partially supported by ITF grant PRP/017/22FX and RGC GRF grant 16218724.